# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v2

@@ III-D Example System: TurtleBot
We use the TurtleBot unicycle model as an illustrative example from the safe planning and control literature [5, 9]. Our planning model is a Dubins car:
@@
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative LQR tracking controller [31] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-2.5,-1.5],p_{y}\in[-1,0]\}$ (see Figures 2 and 3).
@@ IV Proposed Method
In this section, we detail our approach to solving the BRAS problem. First, we convert the planning model to a PWA system to enable our Piecewise Affine Reach-Avoid Computation (PARC). We then adapt the method in [19] to underapproximate the reach set of the PWA planning model without tracking error, to simplify exposition. Next, we propose a novel approach to overapproximate the continuous time avoid set of the PWA system without tracking error (Key result: Theorem 11). To enable incorporating tracking error between the planned and realized trajectories, we adapt the methods in [10, 9] to estimate time-varying tracking error. Finally, we propose PARC to incorporate tracking error and compute the BRAS (Key result: Theorem 17), enabling provably-safe goal-reaching.
@@ IV-A1 Augmented State, Goal, and Obstacles
We perform this reordering and stacking because, as we see later in Lem. 10, the $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so it indexing them as the first states of $\mathbf{x}$ simplifies avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
@@ IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X_{t}$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [32] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
@@
$\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))=\begin{bmatrix}(\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{1:{n}_{w}}\\ \mathbf{k}\\ (\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{{n}_{w}+1:{n}_{p}}\\ \end{bmatrix}$ (36e)
$\mathbf{C}_{i,t}=\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},\quad\mathrm{\textnormal{and}}$ (36f)
$\mathbf{d}_{i,t}=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}.$ (36g)
@@
IV-B Backward Reach Set without Tracking Error
@@
Obtaining expert plans is system-dependent, but is not difficult, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value nonlinear MPC problem, for which solution guarantees exist [33]. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
@@
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$. Then, we can compute the exact $t_{\mathrm{\textnormal{f}}}$-time BRS using the method in [19]:
@@ IV-C PARC: Avoid Set without Tracking Error
Suppose $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in the workspace as in Assum. 2. Also suppose we convert it to a PWA system as in Section IV-A2 and (IV-A2). Then in mode $i$ at time $t$, each state $\mathbf{x}$ of the PWA system obeys
@@
This follows directly from Assum. 2 and (IV-A2). ∎
@@
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{other}}}(t)\end{bmatrix},$ (46)
where $\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\in X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states in $\mathbf{x}(t)$ that are ETI-states, and $\mathbf{x}_{\mathrm{\textnormal{other}}}(t)\in X_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}}$ are the remaining states in $\mathbf{x}(t)$. Thus ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\geq{n}_{w}+{n}_{k}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly. This reordering is only to simplify the exposition.
@@
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and some mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ as in (39). Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{O}_{i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{other}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$, and define the intermediate avoid set $\overline{\Lambda}_{i,t,t}\subset\overline{\Omega}_{t}$ as:
$\overline{\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{other}}},\mathcal{O}_{i}\right)\cap\overline{\Omega}_{t}$ (47)
@@
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [34, 35], or compute the set difference between the reach set and the avoid set [36], and then sample via solving an LP. The reach set and the avoid set are shown in Fig. 3; notice that the avoid set is a nonconvex union of polytopes.
Figure 3: Visualization of the reach set (blue) and the avoid set (red) from Section IV-C for the TurtleBot example (Section III-D and Fig. 2), with an obstacle (light red), extended to all $\theta\in[-\pi,\pi]$. The expert plan used to compute the reach set is shown as a dashed black line; notice that it intersects the obstacle. However, PARC is still able to compute the reach set (blue) and thereby find plans that safely reach the goal set (green).
@@ IV-D Estimating Tracking Error
Our method is based on [10, 37]. We sample $(\mathbf{z}_{0,j},\mathbf{k}_{j})\in\tilde{Z}\times\tilde{K}\subset Z\times K$ for $j=1,\cdots,{n}_{\mathrm{\textnormal{sample}}}$ uniformly, where $\tilde{Z}\times\tilde{K}$ is a hyperrectangle created by partitioning $Z\times K$; we do not index this partition to ease notation. We then have the robot start at $\mathbf{z}_{0,j}$ to track the plans parameterized by initial planning state $\mathbf{p}_{0,j}\triangleq(\mathbf{z}_{0,j})_{1:{n}_{p}}$ and trajectory parameter $\mathbf{k}_{j}$, and compute the error between the planned and realized trajectories in the workspace states (recall that we assume the goal and obstacles are subsets of workspace).
@@
We assume this approach provides an upper bound for feasible tracking error, which can be proven for simple planning models [10, 37].
@@
This assumption is clearly a potential weakness of the method, but we find that it holds in practice because the sampling process produces adversarial $(\mathbf{z}_{0},\mathbf{k})$ pairs. Furthermore, we can refine the tracking error estimate by increasing the fineness of the partition of $Z\times K$ into subsets $\tilde{Z}\times\tilde{K}$ [10].
@@
It is straightforward to express these intervals as H-polytopes [22]. The Cartesian product with empty sets is necessary to augment the states for Minkowski sum and Pontryagin difference operations with the obstacles and the goal.
An example of computed tracking error for a near-hover quadrotor dynamical model is shown in Fig. 8.

### v3

@@ III-D Example System: TurtleBot
We use the TurtleBot unicycle model as an illustrative example from the safe planning and control literature [chen2021fastrack, kousik2020bridging]. Our planning model is a Dubins car:
@@
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative LQR tracking controller [kong2021ilqr] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-1.75,-1.25],p_{y}\in[-0.25,0.25]\}$ (see Figures 2 and 3).
@@ IV Proposed Method
In this section, we detail our approach to solving the BRAS problem. First, we convert the planning model to a PWA system to enable our Piecewise Affine Reach-Avoid Computation (PARC). We then adapt the method in [thomas2006robust] to compute a subset of the reach set of the PWA planning model without tracking error, to simplify exposition. Next, we propose a novel approach to overapproximate the continuous time avoid set of the PWA system without tracking error (Key result: Theorem 13). To enable incorporating tracking error between the planned and realized trajectories, we adapt the methods in [kousik2019safe, kousik2020bridging] to estimate time-varying tracking error. Finally, we propose PARC to incorporate tracking error and compute the BRAS (Key result: Theorem 20), enabling provably-safe goal-reaching.
@@ IV-A1 Augmented State, Goal, and Obstacles
We perform this reordering and stacking because, as we will see later in Lem. 12, the $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so indexing them as the first states of $\mathbf{x}$ will simplify avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
@@ IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X_{t}$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [casselman2009new] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
@@
$\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))=\begin{bmatrix}(\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{1:{n}_{w}}\\ \mathbf{k}\\ (\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{{n}_{w}+1:{n}_{p}}\\ \end{bmatrix}$ (20b)
$\mathbf{C}_{i,t}=\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},\quad\mathrm{\textnormal{and}}$ (20c)
$\mathbf{d}_{i,t}=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}.$ (20d)
@@
From this point onwards, we will be using the PWA system defined by (20) as the planning model instead of (12). That is, the tracking error, BRAS, and reach-avoid guarantees will be computed with respect to the tracking model (14) and the planning model (20). Thus, PARC is concerned with the tracking error between (14) and (20), but not the error induced by affinizing (12) to (20).
Remark 8.
Choosing a larger $\Delta t$ would induce higher mismatch between the tracking model and the planning model, thus increasing the conservativeness of the method. In scenarios with highly dynamic systems or cluttered environments, selecting a smaller $\Delta t$ reduces conservatism, as demonstrated in Fig. 4. As will be shown in the rest of this section, PARC’s reach-avoid guarantees will still be maintained regardless of the choice of $\Delta t$, and its computation time is inversely proportional to $\Delta t$.
IV-B PARC: Reach Set without Tracking Error
@@
Obtaining expert plans is system-dependent, but is not difficult, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value nonlinear MPC problem, for which solution guarantees exist [bonalli2019gusto]. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
@@
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$. Then, we can compute the exact $t_{\mathrm{\textnormal{f}}}$-time BRS using the method in [thomas2006robust]:
@@
Traditionally, methods such as the reachableSet function in MPT3 [herceg2013multi] compute the full BRS of the PWA system instead of that of only a particular mode sequence. While computing the full BRS would return all possible reachable states and do not require knowledge of an expert plan, the computation time grows exponentially with the number of BRS steps and quickly becomes intractable due to the large number of polytopes, as shown in Table I. Instead, Proposition 10 boasts a much more reasonable computation time by isolating the affine dynamics of a single mode sequence and keeping only one polytope regardless of the number of BRS steps.
@@
Computation Time (s) Number of Polytopes
BRS Steps MPT3 Prop. 10 MPT3 Prop. 10
1 0.03 0.0015 16 1
2 0.27 0.0022 100 1
3 1.45 0.0038 484 1
4 6.63 0.0036 2116 1
5 29.46 0.0027 8836 1
26 timeout 0.0155 1
TABLE I: Computation time and number of polytopes for 1, 2, 3, 4, 5, and 26-step BRS for the TurtleBot example (Section III-D) using the reachableSet function in MPT3 [herceg2013multi] and Proposition 10. Since MPT3 computes the full BRS of the PWA system, the computation time quickly grows intractable due to the large number of polytopes. In contrast, Proposition 10 maintains a reasonable computation time by keeping only one polytope at each step, characterized by the affine dynamics corresponding to the mode sequence of the expert plan.
@@ IV-C PARC: Avoid Set without Tracking Error
Suppose $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ is an ETI system in the workspace as in Assum. 2. Also suppose we convert it to a PWA system as in Section IV-A2 and (20). Then in mode $i$ at time $t$, each state $\mathbf{x}$ of the PWA system obeys
@@
This follows directly from Assum. 2 and (20). ∎
@@
$\mathbf{x}(t)$ $=\begin{bmatrix}\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\\ \mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\end{bmatrix},$ (28)
where $\mathbf{x}_{\mathrm{\textnormal{ETI}}}(t)\in X_{\mathrm{\textnormal{ETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}$ are the first ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}$ states in $\mathbf{x}(t)$ that are ETI states, and $\mathbf{x}_{\mathrm{\textnormal{nonETI}}}(t)\in X_{\mathrm{\textnormal{nonETI}}}\subset\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}}$ are the remaining states in $\mathbf{x}(t)$. Thus ${{n}_{x}}_{\mathrm{\textnormal{ETI}}}\geq{n}_{w}+{n}_{k}$. Note that all sets and matrices including $\mathcal{O}$, and $P_{\mathrm{\textnormal{G}}}$ are relabeled accordingly. This reordering is only to simplify the exposition.
@@
Consider the obstacle $\mathcal{O}_{i}$, some time $t\in\{0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$, and some mode sequence $S=(s_{0},\cdots,s_{t},s_{t+\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t})$. Suppose we compute the $(t_{\mathrm{\textnormal{f}}}-t)$-time BRS $\overline{\Omega}_{t}$ as in (23). Let $A=\mathcal{B}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(\mathcal{O}_{i}\right)\times\mathbb{R}^{{{n}_{x}}_{\mathrm{\textnormal{nonETI}}}},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)$, and define the intermediate avoid set $\overline{\Lambda}_{i,t,t}\subset\overline{\Omega}_{t}$ as:
$\overline{\Lambda}_{i,t,t}=\mathrm{\textnormal{conv}}\!\left(\mathrm{\textnormal{proj}}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\!\left(A\right)\times X_{\mathrm{\textnormal{nonETI}}},\mathcal{O}_{i}\right)\cap\overline{\Omega}_{t}$ (29)
@@
Remark 14.
We note that PARC can be readily extended to moving obstacles by replacing $\mathcal{O}_{i}$ in (29) with the over-approximated swept volume of the obstacle between $t$ and $\Delta t$, assuming its motion is known or can be over-approximated. This strategy has shown to be effective on other set-based planner-tracker frameworks [vaskov2019towards, vaskov2019not]. For ease of exposition, we consider only static obstacles in this paper, as many of the methods we are comparing PARC against do not handle moving obstacles [chen2021fastrack, dawson2022safe, verginis2022kdf], and reach-avoid for static obstacles in near-danger setups is already a difficult problem for existing literature.
@@
To sample from within the reach set but not from the avoid set, one can either rejection-sample, sample via random-walk algorithms [kannan2009random, chen2018fast], or compute the set difference between the reach set and the avoid set [baotic2009polytopic], and then sample via solving an LP. An example of the reach set and the avoid set are shown in Fig. 3; notice that the avoid set is a nonconvex union of polytopes.
Figure 3: (Left) Visualization of the reach set (green) and the avoid set (red) computation from Section IV-C for the TurtleBot example (Section III-D), sliced in the first three dimensions with respect to the initial condition $\mathbf{p}_{0}=[-3.5,-0.5,\pi/5]^{\intercal}$. The black star denotes the expert plan with $\mathbf{k}=[-0.318,0.9]^{\intercal}$ and the blue star denotes a safe plan with $\mathbf{k}=[-0.320,1.1]^{\intercal}$. (Right) Illustration of the resulting trajectory in $p_{x}$ and $p_{y}$, with the expert plan shown as a dashed black line and the safe plan shown as a dashed blue line. While the expert plan collides with the obstacle (red), PARC is still able to compute a safe plan with the same mode sequence that safely reach the goal set (green).
Figure 4: Visualization of the reach set (green), the avoid set (red), and a grid-based over-approximation of the true avoid set (yellow) for the TurtleBot example (Section III-D) for different values of the timestep $\Delta t$, with the same problem setup and slicing as Fig. 3. The grid-based over-approximation was obtained by dividing the reach set into a $50\times 50$ grid, and checking if the trajectories from each grid collides with the obstacle using FRS computation [herceg2013multi] and Proposition 11. Both the reach set and the numerical overapproximation error gets smaller as $\Delta t$ decreases.
Disregarding tracking error and perception uncertainties, which all planning and control methods suffer from, PARC’s only source of conservativeness (from computing the BRAS with respect to the expert plan’s mode sequence) comes from the convex hull approximation for linearly interpolated trajectories and the projection operation to account for non-ETI states. We illustrate the extent of this overapproximation in Fig. 4. Note that the amount of overapproximation depends heavily on the system, the problem setup, and the expert plan, but can generally be alleviated by choosing smaller timesteps at the expense of computation time.
@@ IV-D Estimating Tracking Error
Our method is based on [kousik2019safe, shao2021reachability]. We sample $(\mathbf{z}_{0,j},\mathbf{k}_{j})\in\tilde{Z}\times\tilde{K}\subset Z\times K$ for $j=1,\cdots,{n}_{\mathrm{\textnormal{sample}}}$ uniformly, where $\tilde{Z}\times\tilde{K}$ is a hyperrectangle created by partitioning $Z\times K$; we do not index this partition to ease notation. We then have the robot start at $\mathbf{z}_{0,j}$ to track the plans parameterized by initial planning state $\mathbf{p}_{0,j}\triangleq(\mathbf{z}_{0,j})_{1:{n}_{p}}$ and trajectory parameter $\mathbf{k}_{j}$, and compute the error between the planned and realized trajectories in the workspace states (recall that we assume the goal and obstacles are subsets of workspace).
@@
We assume this approach provides an upper bound for feasible tracking error, which can be proven for simple planning models [kousik2019safe, shao2021reachability].
@@
This assumption is clearly a potential weakness of the method, but we find that it holds in practice because the sampling process produces adversarial $(\mathbf{z}_{0},\mathbf{k})$ pairs. Furthermore, we can refine the tracking error estimate by increasing the fineness of the partition of $Z\times K$ into subsets $\tilde{Z}\times\tilde{K}$ [kousik2019safe]. Note there exist alternative methods to collect or construct tracking error functions in the literature [verginis2022kdf, chen2021fastrack]. If desired, they could be swapped out with (33) and (34) and PARC would provide guarantees with respect to the new tracking error functions instead. We demonstrate PARC’s modularity to the tracking error model in Section V-D.
@@
It is straightforward to express these intervals as H-polytopes [herceg2013multi]. The Cartesian product with singletons of zeros is necessary to augment the states for Minkowski sum and Pontryagin difference operations with the obstacles and the goal.
We note that collection of the tracking error only requires forward simulation or rollouts of the system and knowledge of the input and output states, so our approach could still apply if the tracking model is a “black box” by adapting methods such as [selim2022safe, alanwar2023data]. In this paper, we consider “white-box” dynamics in our experiments to treat them as a control condition to compare against the literature. While the forward simulation and collision checking of trajectories may sometimes be computationally expensive, we consider the tradeoff acceptable as the tracking error of PARC is expected to be collected offline. An example of computed tracking error for a near-hover quadrotor dynamical model is shown in Fig. 9.